OpenStudy (anonymous):

A 10-foot ladder leans against a wall with its foot braced 3 feet from wall’s base. How far up the wall does the ladder reach?

OpenStudy (anonymous):

\[\sqrt{10^2-3^2}=\sqrt{91}=9.54 \]

OpenStudy (anonymous):

ok thats what i got 9.54 feet up the wall. i just wanted to make sure thanks

OpenStudy (anonymous):

You're welcome.